# Primitive-Root Determinant Densities over Prime Fields and Implications for PRIM-LWE

### 30 Mar 2026 v3

@@ 6 Explicit Lower Bounds for Primes of Cryptographic Interest
For the general bound, since every prime divisor of q1q-1 is at most q1q-1,
$\frac{\varphi(q-1)}{q-1}=\prod_{r\mid(q-1)}\Bigl(1-\frac{1}{r}\Bigr)\;\geq\;\prod_{\begin{subarray}{c}r\leq q-1\\ r\text{ prime}\end{subarray}}\Bigl(1-\frac{1}{r}\Bigr),$
where the inequality holds because the right-hand side includes additional factors in $(0,1)$.
By Dusart’s explicit product estimate [20, Theorem 5.9], which we use conservatively with threshold $x\geq 2\,278\,383$,
@@
Since $q>2^{30}$, we have q1>2 278 383q-1>2\,278\,383 and $\log(q-1)>20$, so the correction factor lies in $(1-2.5\times 10^{-5},\;1+2.5\times 10^{-5})$.
@@
$\prod_{j\geq 1}\Bigl(1-\frac{1}{q^{j}}\Bigr)\geq 1-\sum_{j\geq 1}\frac{1}{q^{j}}=1-\frac{1}{q-1}>0.9999.$
@@
$c(q)=\frac{\varphi(q-1)}{q-1}\prod_{j\geq 1}\Bigl(1-\frac{1}{q^{j}}\Bigr)\;>\;\Bigl(1-\frac{0.2}{\log^{3}(q-1)}\Bigr)\cdot 0.9999\cdot\frac{e^{-\gamma}}{\log(q-1)}\;>\;\frac{0.56}{\log q},$
using $e^{-\gamma}>0.5614$, $\log(q-1)<\log q$, and $0.2/\log^{3}(q-1)<2.5\times 10^{-5}$ for $q>2^{30}$. ∎
The bound $c(q)>0.56/\!\log q$ of Corollary 10 is a simple, fully explicit estimate obtained by taking the Mertens product over all primes up to q1q-1. The sharper universal pointwise bound $c(q)\gg 1/\!\log\log q$ of Corollary 9 shows that $1/\!\log\log q$, not $1/\!\log q$, is the correct worst-case scale for the rejection-sampling overhead.
@@
By the shifted-prime Erdős–Kac theorem (Barban [24]; see also Halberstam [25] and Elliott [26, Ch. 12]), $(\omega(q-1)-\log\log x)/\sqrt{\log\log x}$ converges in distribution to the standard normal as $x\to\infty$ over primes $q\leq x$. In particular, for every fixed $\delta>0$, almost all primes $q\leq x$ satisfy $\omega(q-1)\leq(1+\delta)\log\log x$.
Remark 4 (Typical size of $\omega(q-1)$).
The shifted-prime Erdős–Kac theorem explains why bounds depending only on $\omega(q-1)$ are relevant in cryptographic parameter ranges: for $q\approx 2^{128}$, one has $\log\log q\approx 4.5$, so $\omega(q-1)\leq 10$ holds for the vast majority of such primes.
Scheme $q$ factorization of q1q-1 general lower bound for $c(q)$ exact-factorization lower bound
ML-KEM $3329$ $2^{8}\cdot 13$ $B(2)=P_{3}/3\approx 0.1867$ $\dfrac{6}{13}\Bigl(1-\dfrac{1}{3328}\Bigr)\approx 0.4614$
ML-DSA $8380417$ $2^{13}\cdot 3\cdot 11\cdot 31$ $B(4)=8P_{3}/35\approx 0.1280$ $\dfrac{100}{341}\Bigl(1-\dfrac{1}{8380416}\Bigr)\approx 0.2933$
@@
Table 1: The prime moduli appearing in the current NIST standards for ML-KEM [22] and ML-DSA [23]. The fourth column gives the general $\omega(q-1)$-based lower bound for $c(q)$ from Theorem 11, while the fifth gives the sharper lower bound obtained from the exact factorization of q1q-1 together with $P(q)\geq 1-\frac{1}{q-1}$. The exact-factorization bounds imply dimension-uniform expected rejection-sampling overheads at most $2.17$ and $3.42$, respectively.
For the two prime moduli appearing, across all parameter sets, in the current NIST standards for ML-KEM and ML-DSA [22, 23], Table 1 records both the general $\omega(q-1)$-based bounds and the sharper exact-factorization bounds. The exact factorizations give $\varphi(3328)/3328=6/13$ and $\varphi(8380416)/8380416=100/341$. Together with $P(q)\geq 1-\frac{1}{q-1}$, one obtains
$c(3329)\geq\frac{6}{13}\Bigl(1-\frac{1}{3328}\Bigr)\approx 0.4614,\qquad c(8380417)\geq\frac{100}{341}\Bigl(1-\frac{1}{8380416}\Bigr)\approx 0.2933.$
A short tail-bound argument (taking $J=2$ partial products of $P(q)$ and using the estimate $P_{J}(q)-P(q)\leq q^{-J}/(q-1)$) confirms that these lower bounds are tight to many decimal places:
$c(3329)=0.4613997781\ldots,\qquad c(8380417)=0.2932550969\ldots.$
The corresponding dimension-uniform expected rejection-sampling overheads $1/c(q)$ are approximately $2.17$ and $3.41$. For each fixed dimension $n$, the exact overheads $1/c_{n}(q)$ are no larger.
Context $q$ factorization of q1q-1 $\omega(q-1)$ $1/c(q)$
ML-KEM (FIPS 203) $3329$ $^{8}\cdot 13 22$ $2.17$
ML-DSA (FIPS 204) $8380417$ $2^{13}\cdot 3\cdot 11\cdot 31$ 44 $3.41$
NewHope / FALCON [27, 28] $12289$ $^{12}\cdot 3 22$ $3.00$
BFV/BGV FHE (small) [29, 30, 31] $786433$ $^{18}\cdot 3 22$ $3.00$
SEAL BFV default [32] $132120577$ $2^{21}\cdot 3^{2}\cdot 7$ 33 $3.50$
BFV/BGV FHE (medium) [29, 30, 31] $2013265921$ $2^{27}\cdot 3\cdot 5$ 33 $3.75$
NTTRU / BLISS [33, 34] $7681$ $2^{9}\cdot 3\cdot 5$ 33 $3.75$
64-bit NTT prime [35] $2^{64}-2^{32}+1$ $2^{32}\cdot 3\cdot 5\cdot 17\cdot 257\cdot 65537$ 66 $4.00$
Fermat prime ($2^{16}+1$) $65537$ $2^{16}$ 11 $2.00$
@@
Table 2: Representative NTT-friendly primes from lattice-based cryptography and FHE implementations, together with $\omega(q-1)$ and the dimension-uniform expected rejection-sampling overhead $1/c(q)$. The displayed overheads are rounded to two decimal places; the exact values differ by less than $10^{-3}$ due to the Euler product $P(q)$. All overheads are at most $4.00$. The first two rows reproduce the NIST entries from Table 1.
Table 2 collects nine representative moduli drawn from current lattice-based and FHE implementations. Every modulus in the table satisfies $\omega(q-1)\leq 6$ and incurs a dimension-uniform rejection-sampling overhead $1/c(q)\leq 4.00$. The table substantiates the claim that the primes typically encountered in practice have controlled factorization structure, which yields a good lower bound on $c(q)$.
@@
For the 64-bit NTT prime $q=2^{64}-2^{32}+1$ (sometimes called the “Goldilocks prime” in the zero-knowledge community, not to be confused with the Ed448-Goldilocks prime $2^{448}-2^{224}-1$ [36]), used as an NTT-friendly ambient field in implementation-oriented lattice cryptography [35], one has $\omega(q-1)=6$ and $\varphi(q-1)/(q-1)\approx 0.2500$. Since $P(q)>1-10^{-19}$, one obtains $c(q)\approx 0.25$, so in every dimension $n$ at least a fraction $\approx 0.25$ of random $n\times n$ matrices have a primitive-root determinant, and the dimension-uniform security-reduction overhead is about $4$.
7 Implications for PRIM-LWE
For a fixed prime modulus $p$, [1, Theorem 10] shows that decision-prim-lwe is at least as hard as decision-LWE, up to an $O(n^{2})$ factor. Separately, [1, Proposition 5] isolates a positive lower bound, uniform in the matrix dimension, for the probability that a random secret matrix has primitive-root determinant. In our notation, that dimension-uniform sampling constant is
$c=\inf_{n\geq 1}c_{n}(p).$
As shown in Section 2, $c_{n}(p)$ is decreasing in $n$ and $\inf_{n\geq 1}c_{n}(p)=c(p)$, so $c(p)$ is the dimension-uniform sampling term entering the reduction. The present note concerns this quantity $c(p)$; it does not alter the separate $O(n^{2})$ factor from [1, Theorem 10].
Theorem 3 shows that no uniform lower bound on $c(p)$ holds across all primes, meaning the reduction tightness degrades for primes $p$ whose predecessor p1p-1 is divisible by many small primes. More precisely, let $p^{*}(k)$ denote the least prime satisfying $p^{*}(k)\equiv 1\pmod{N_{k}}$, as in the proof of Proposition 1. Then, for the absolute constants $C_{0},L$ from Theorem 4,
$N_{k}<p^{*}(k)\leq C_{0}N_{k}^{\,L}.$
The left inequality holds because $p^{*}(k)=1+jN_{k}$ with $j\geq 1$, while the right inequality is exactly Linnik’s theorem. Therefore
$\log p^{*}(k)\asymp\log N_{k},\qquad\log N_{k}\sim p_{k}$
by Lemma 2, and hence
$\log\log p^{*}(k)\asymp\log p_{k}.$
On the other hand, the proof of Proposition 1 gives
$c\bigl(p^{*}(k)\bigr)\leq\prod_{i=1}^{k}\Bigl(1-\frac{1}{p_{i}}\Bigr)\sim\frac{e^{-\gamma}}{\log p_{k}}.$
Thus
$c\bigl(p^{*}(k)\bigr)\ll\frac{1}{\log\log p^{*}(k)},$
so the construction yields an infinite sequence of primes $p=p^{*}(k)$ for which
$c(p)=O\!\left(\frac{1}{\log\log p}\right).$
The dimension-uniform upper bound on the rejection-sampling overhead, namely $1/c(p)$, can therefore grow like
$\Omega(\log\log p).$
In fixed dimension $n$, the exact expected number of trials is $1/c_{n}(p)\leq 1/c(p)$. By [1, Theorem 10], the reduction from decision-LWE to decision-prim-lwe introduces an advantage loss of $O(n^{2})\cdot(1/c)$, where $c=\inf_{n\geq 1}c_{n}(p)=c(p)$ is the dimension-uniform sampling constant. Our results therefore show that the $1/c(p)$ factor in this bound can grow like $\Omega(\log\log p)$ along a suitable subsequence of primes.
However, this is a modulus-dependent artifact rather than a structural obstruction. For any fixed prime $p$, the constant $c(p)>0$ and the reduction remains valid. Moreover, the explicit bounds of Section 6 show that the specific moduli surveyed in Table 2—which cover the NIST standards and several widely deployed FHE primes—all have $c(q)$ concretely bounded away from zero.
Corollary 3 gives $\limsup c(p)=1/2$ unconditionally, so there is no shortage of primes for which $c(p)$ is close to its maximal value. As a concrete illustration, for safe primes $q=2r+1$ with $r$ an odd prime and $q\geq 7$, one has
$c(q)=\frac{r-1}{2r}\prod_{j\geq 1}\left(1-q^{-j}\right)\geq P_{3}\,\frac{r-1}{2r}.$
The prefactor (r1)/(2r)=1/21/(q1)(r-1)/(2r)=1/2-1/(q-1) is already close to $1/2$ when $q$ is large, and the Euler product $\prod_{j\geq 1}(1-q^{-j})$ tends to $1$ as $q\to\infty$, so $c(q)$ approaches $1/2$ for large safe primes.
More generally, Corollary 4 shows that for every fixed $\tau\in(0,1/2)$, both sets of primes
$\{p:c(p)\leq\tau\}\qquad\text{and}\qquad\{p:c(p)>\tau\}$
have positive relative density, namely $G(\tau)$ and $1-G(\tau)$. Moreover, $G(\tau)\to 0$ as $\tau\downarrow 0$. Thus moduli yielding a very loose constant-factor sampling term are rare only in the threshold-asymptotic sense $\tau\downarrow 0$, not absent for any fixed threshold $\tau<1/2$.
We note that the primes typically used in lattice-based cryptography are chosen for NTT compatibility. Concretely, this means $q\equiv 1\pmod{2^{t}}$ for moderate $t$. Mathematically, this condition controls only the 2-adic part of q1q-1. It does not by itself prevent q1q-1 from having many odd prime factors, and NTT-friendliness alone does not imply a good lower bound on $c(q)$. What our theorems actually use is the controlled factorization structure of q1q-1, quantified by $\omega(q-1)$ or by the exact factorization. The NTT condition is harmless with respect to the 2-part, since the large power of $2$ contributes only the single Euler factor $(1-\tfrac{1}{2})$ to $\varphi(q-1)/(q-1)$. Among the implementation moduli surveyed in Table 2, every prime satisfies $\omega(q-1)\leq 6$. It is this controlled factorization that yields the good lower bounds on $c(q)$ recorded in Remark 5, and Tables 1 and 2. Thus our results do not by themselves indicate a practical weakness; they quantify how modulus choice affects the reduction loss in the prim-lwe framework.
Affine determinant programs (ADPs) [37] provide a natural external determinant-based connection to prim-lwe. In the ADP obfuscation framework, a Kilian-style randomization step left- and right-multiplies by invertible matrices $R,S$ with $\det(R)\det(S)=1$ to preserve the determinant polynomial. Because the ambient sampling space is $\mathrm{GL}_{n}(\mathbb{F}_{q})$ rather than all of $M_{n}(\mathbb{F}_{q})$, the relevant primitive-root fraction reduces to $\varphi(q-1)/(q-1)$: the matrix-count Euler product $\prod_{j=1}^{n}(1-q^{-j})$ is already absorbed into the $\mathrm{GL}_{n}$ normalization. Since the determinant map
$\det:\mathrm{GL}_{n}(\mathbb{F}_{q})\longrightarrow\mathbb{F}_{q}^{\times}$
is a surjective homomorphism, all fibers have the same cardinality, and therefore the exact fraction of invertible matrices with primitive-root determinant is
$\frac{\varphi(q-1)}{q-1}.$
Hence, if $\omega(q-1)\leq 10$, then
$\frac{\varphi(q-1)}{q-1}\;\geq\;\prod_{i=1}^{10}\Bigl(1-\frac{1}{p_{i}}\Bigr)\approx 0.15795,$
so rejection sampling inside $\mathrm{GL}_{n}(\mathbb{F}_{q})$ succeeds in fewer than $6.4$ trials on average. If instead one samples from all matrices $M_{n}(\mathbb{F}_{q})$, then the exact finite-$n$ primitive-root density is $c_{n}(q)$, whose limit as $n\to\infty$ is $c(q)$. While our results do not establish a security advantage for imposing primitive-root determinant constraints, they confirm that the associated sampling overhead is mild whenever q1q-1 has controlled factorization.
@@ 8 Concluding Remarks
We have resolved the open question posed in [1, Remark 3(ii)] in the affirmative: $\inf_{p\text{ prime}}c(p)=0$, and in fact
$\liminf_{\begin{subarray}{c}p\to\infty\\ p\text{ prime}\end{subarray}}c(p)=0.$
The proof is short and unconditional, requiring only Dirichlet’s theorem and Mertens’ product formula. We have further established several results that describe the finer structure of the density $c(p)$:
  1. 1.
    The minimum of $c(p)$ up to $x$ has the sharp order
    In particular, the decay to zero is slow and quantified precisely.
  2. 2.
    The quantity $c(p)$ has a continuous limiting distribution over the primes whose support is exactly $0,\tfrac{1}{2}]$; equivalently, the distribution function is strictly increasing on $0,\tfrac{1}{2}]$. This full-support conclusion appears to be new. The law coincides with the classical shifted-prime distribution of $\varphi(p-1)/(p-1)$, is purely singular, and, by Proposition 4, the limiting measure has Hausdorff dimension 0. The moments of $G$ are given by convergent Euler products (Theorem 8), and the Mellin transform $\mathbb{E}[X^{s}]$ extends analytically to $\{\operatorname{Re}(s)>0\}$ (Theorem 10). Pure singularity was already known in the integer case from Erdős [5]; Tenenbaum [3, p. 263] noted that it carries over to shifted primes.
  3. 3.
    For the implementation moduli surveyed in Table 2—including the current NIST-standardized moduli and several widely used FHE primes—the density $c(q)$ admits an explicit positive lower bound depending only on $\omega(q-1)$. Accordingly, the prim-lwe reduction has a finite, explicitly computable constant-factor sampling term for every fixed prime modulus, and this term is small in the concrete parameter regimes treated here.
On the cryptographic side, the rejection-sampling overhead $1/c(p)$ can grow like $\Omega(\log\log p)$ along a suitable subsequence of primes, but as discussed in Section 7, this represents a quantitative sensitivity to modulus selection rather than a structural obstruction. Table 2 confirms that every implementation modulus we examined has $\omega(q-1)\leq 6$ and $1/c(q)\leq 4.00$, and the sharper universal pointwise bound of Corollary 9 gives $1/c(q)=O(\log\log q)$.
Proposition 7 and Corollary 7 determine the right-endpoint behavior of $G$:
$1-G\!\left(\frac{1}{2}-\varepsilon\right)\sim\frac{\kappa}{\log(1/\varepsilon)}\qquad(\varepsilon\downarrow 0),$
so $G$ is not Hölder of any positive order at $1/2$. Proposition 4 shows that the limiting measure has Hausdorff dimension 0. Equivalently, the limiting PRIM-LWE law is a continuous, purely singular, full-support measure on $0,\tfrac{1}{2}]$ that is concentrated on a Borel set of Hausdorff dimension 0. A natural remaining question is therefore to determine a global modulus of continuity of $G$. The squared modulus of the characteristic function admits an explicit Euler product (Proposition 5), but whether

### 21 Apr 2026 v4

@@ 7 Explicit Lower Bounds via the Number of Distinct Prime Factors
For the general bound, since every prime divisor of p1p-1 is at most p1p-1,
$\frac{\varphi(p-1)}{p-1}=\prod_{r\mid(p-1)}\Bigl(1-\frac{1}{r}\Bigr)\;\geq\;\prod_{\begin{subarray}{c}r\leq p-1\\ r\text{ prime}\end{subarray}}\Bigl(1-\frac{1}{r}\Bigr).$
By Dusart’s explicit product estimate [12, Theorem 5.9], which we use with threshold $x\geq 2\,278\,382$ (noting that $2\,278\,382<2^{22}\ll 2^{30}$, so the threshold is satisfied for all primes $p>2^{30}$),
@@
Since $p>2^{30}$, we have p1>2 278 382p-1>2\,278\,382 and $\log(p-1)>20$, so the correction factor lies in $(1-2.5\times 10^{-5},\;1+2.5\times 10^{-5})$.
@@
$\prod_{j\geq 1}\Bigl(1-\frac{1}{p^{j}}\Bigr)\geq 1-\sum_{j\geq 1}\frac{1}{p^{j}}=1-\frac{1}{p-1}>0.9999.$
@@
$c(p)=\frac{\varphi(p-1)}{p-1}\prod_{j\geq 1}\Bigl(1-\frac{1}{p^{j}}\Bigr)\;>\;\Bigl(1-\frac{0.2}{\log^{3}(p-1)}\Bigr)\cdot 0.9999\cdot\frac{e^{-\gamma}}{\log(p-1)}\;>\;\frac{0.56}{\log p},$
using $e^{-\gamma}>0.5614$, $\log(p-1)<\log p$, and $0.2/\log^{3}(p-1)<2.5\times 10^{-5}$ for $p>2^{30}$. Equivalently, for $p>2^{30}$,
$\frac{1}{c(p)}\;<\;\frac{\log p}{0.56}\;<\;1.79\log p.$
Remark 7.
@@
By the shifted-prime Erdős–Kac theorem (Barban [3]; see also Halberstam [20] and Elliott [13, Ch. 12]), extending the classical result of Erdős and Kac [15], $(\omega(p-1)-\log\log x)/\sqrt{\log\log x}$ converges in distribution to the standard normal over primes $p\leq x$. For $p\approx 2^{128}$, $\log\log p\approx 4.5$, so $\omega(p-1)\leq 10$ holds for the vast majority of such primes.
Standard $p$ generic $B(\omega(p-1))$ bound sharper factorization bound
ML-KEM $3329$ $B(2)=P_{3}/3\approx 0.1867$ $\approx 0.4614$
ML-DSA $8380417$ $B(4)=8P_{3}/35\approx 0.1280$ $\approx 0.2933$
@@
Table 2: The prime moduli appearing in FIPS 203 and FIPS 204 [31, 32]. The third column gives the generic bound $B(\omega(p-1))$ from Theorem 9, depending only on the number of distinct prime factors of p1p-1, while the fourth gives the sharper lower bound obtained from the exact factorization of p1p-1 together with $P(p)\geq 1-\frac{1}{p-1}$.
For the NIST-standardized moduli [31, 32], the exact factorizations $\varphi(3328)/3328=6/13$ and $\varphi(8380416)/8380416=100/341$ together with $P(p)\geq 1-1/(p-1)$ give $c(3329)\geq 0.4613$ and $c(8380417)\geq 0.2932$. The certified evaluation of Proposition 8 yields $c(3329)=0.461399\ldots$ and $c(8380417)=0.293255\ldots$, so the sharper factorization bounds agree with the exact values to $10$ and $21$ decimal places, respectively; the gap is $O(p^{-3})$, since replacing $P(p)$ by 11/(p1)1-1/(p-1) incurs an error of order $p^{-1}\cdot p^{-2}$ from the leading cross-term in the Euler product expansion.
Context $p$ $\omega(p-1)$ $1/c(p)$
ML-KEM (FIPS 203) $3329$ 22 $2.17$
ML-DSA (FIPS 204) $8380417$ 44 $3.41$
NewHope / FALCON [2, 19] $12289$ 22 $3.00$
BFV/BGV FHE (small) [5, 18, 6] $786433$ 22 $3.00$
BFV/BGV FHE (medium) [5, 18, 6] $2013265921$ 33 $3.75$
NTTRU / BLISS [27, 11] $7681$ 33 $3.75$
64-bit NTT prime [37] $2^{64}-2^{32}+1$ 66 $4.00$
Fermat prime ($2^{16}+1$) $65537$ 11 $2.00$
@@
Table 3: Representative NTT-friendly primes from lattice-based cryptography and FHE implementations, together with $\omega(p-1)$ and the dimension-uniform expected rejection-sampling overhead $1/c(p)$. The displayed overheads are rounded to two decimal places; the exact values differ by less than $3\times 10^{-3}$.
@@
For the 64-bit NTT prime $p=2^{64}-2^{32}+1$ (sometimes called the “Goldilocks prime”; not to be confused with the Ed448-Goldilocks prime [21]), used in lattice cryptography [37], one has $\omega(p-1)=6$, $\varphi(p-1)/(p-1)\approx 0.2500$, $P(p)>1-10^{-19}$, and $1/c(p)\approx 4$.
Proposition 8 (Certified evaluation of $c(p)$).
Given a prime $p\geq 3$ and the complete factorization $p-1=r_{1}^{e_{1}}\cdots r_{k}^{e_{k}}$, the following procedure computes $c(p)$ to $D$ correct decimal digits using $O(k+D\log 10/\!\log p)$ rational-arithmetic operations in the unit-cost rational-arithmetic model:
  1. (i)
    Compute the totient ratio $\varphi(p-1)/(p-1)=\prod_{i=1}^{k}(1-1/r_{i})$, which is exact as a rational number.
  2. (ii)
    Set the truncation depth $J:=\lceil D\log 10/\!\log p\rceil$ and compute the partial Euler product $P_{J}(p):=\prod_{j=1}^{J}(1-p^{-j})$, again exact in rational arithmetic.
  3. (iii)
    Return the certified interval
    $\frac{\varphi(p-1)}{p-1}\!\left(P_{J}(p)-\frac{p^{-J}}{p-1}\right)\;\leq\;c(p)\;\leq\;\frac{\varphi(p-1)}{p-1}\,P_{J}(p).$
The width of this certified interval is at most $p^{-J}/(2(p-1))$. Moreover, in step (ii) the exact rational $P_{J}(p)$ has denominator $p^{J(J+1)/2}$, so the intermediate numerators and denominators have $O(J^{2}\log p)$ bits.
Proof.
Steps (i) and (ii) use k1k-1 and J1J-1 multiplications respectively. The certified bound in (iii) follows from the identity $P(p)=P_{J}(p)\prod_{j>J}(1-p^{-j})$ together with $\prod_{j>J}(1-p^{-j})\geq 1-\sum_{j>J}p^{-j}=1-p^{-J}/(p-1)$ and $P(p)\leq P_{J}(p)$, giving $0\leq P_{J}(p)-P(p)\leq p^{-J}/(p-1)$. The interval width bound uses $\varphi(p-1)/(p-1)\leq 1/2$. ∎
@@ 8 Concluding Remarks
This article studies the PRIM-LWE density constant $c(p)$ unconditionally from extremal, distributional, and computational viewpoints. We proved that $\inf_{p\text{ prime}}c(p)=0$ and that $\min_{p\leq x}c(p)\asymp 1/\log\log x$, so the worst-case rejection-sampling overhead grows only on a doubly logarithmic scale. We also showed that $c(p)$ admits a continuous, purely singular limiting distribution with support exactly $0,\tfrac{1}{2}]$ and Hausdorff dimension 0; in particular, the limiting measure is supported on a metrically negligible subset of $0,\tfrac{1}{2}]$ despite having full topological support. Finally, the explicit $\omega(p-1)$-based bounds and the certified evaluation algorithm of Proposition 8 make these asymptotic results effective for concrete cryptographic moduli.
Several questions remain open. The squared modulus of the characteristic function admits an explicit Euler product (Proposition 5), but whether